Can someone please help me with this word problem! Thanks a lot.
If $3000 is invested in an account that pays interest compounded continuously, how long will it take to grow $6000 at 7%?

Answer by nycsharkman(136)

If $3000 is invested in an account that pays interest compounded continuously, how long will it take to grow $6000 at 7%?
The formula need is:
A = Pe^rt, where A = amount, P = principal, r = rate (our percent) and t = time
A = 6000
r = 0.07
t = what we are looking for
P = 3000
We now plug and chug.
6000 = 3000e^(0.07)(t)
First divide both sides by 3000.
6000/3000 = e^(0.07)(t)
2 = e^(0.07)(t)
NEXT: Rewrite as a log.
0.07t = In2
Divide both sides by 0.07 to fnd t.
t = In2/0.07
t = 9.902102579
Round 9.902102579 to the nearest ones place and we get 10.
The answer is 10 years.